# Beatty solutions of almost Golomb equations

### 16 Apr 2026 v2

@@ Abstract
The Golomb equation of order $r$ is the implicit functional equation
@@
for nondecreasing sequences of positive integers. Its earliest solution, the almost Golomb sequence of order $r$, is $r$-regular in the sense of Allouche and Shallit and has oscillating ratio $a(n)/n$. We prove that for every $r\geq 2$ that is not an even perfect square, the equation admits a second monotone solution given by an inhomogeneous Beatty sequence.
@@
Keywords. Beatty sequence, almost Golomb sequence, functional equation, Sturmian word, Wythoff game.
@@ 1.2 The greedy solution
The greedy (earliest) solution of (1) is studied in detail in the companion paper [2]. It is unique, $r$-regular, and has explicit $r$-adic denesting formulas. For $r=2$ (A394217 in [8]), it begins
@@ 1.3 A second solution
The two solutions agree on $[1,11]$ and diverge at $n=12$, where the greedy repeats the value $8$ (run of length $2$), while the Beatty moves to $9$ (run of length $1$). This is possible because the constraint at $n=12$ applies at position $f(12)=a(12)+a(11)$, which is $16$ for the greedy and $17$ for the Beatty. Different values of $a$ shift the constrained positions, and both are self-consistent.
@@
Figure 1 shows both solutions over the range $n\leq 512$. The greedy $a_{\mathrm{G}}$ crosses the Beatty $a_{\mathrm{B}}$ repeatedly: the residual $a_{\mathrm{B}}(n)-n/\!\sqrt{2}$ is bounded (it is the Sturmian remainder $\{(n{+}1)/\!\sqrt{2}\}$), whereas $a_{\mathrm{G}}(n)-n/\!\sqrt{2}$ grows without bound because $a_{\mathrm{G}}(n)/n$ oscillates in $[2/3,3/4]$ rather than converging to $1/\!\sqrt{2}$. Figure 2 below illustrates the difference between $a_{\mathrm{G}}$ and $a_{\mathrm{B}}$.
Figure 1: The greedy solution $a_{\mathrm{G}}(n)$ and the canonical Beatty $a_{\mathrm{B}}(n)$ for $1\leq n\leq 512$.
Figure 2: The normalised difference $(a_{\mathrm{G}}(n)-a_{\mathrm{B}}(n))/n$ for $1\leq n\leq 512$.
@@
When $r=(2m)^{2}$, the left-hand side equals $n+1$ instead of $n$ for all $n\geq r$ (Theorem 12).
The proof combines three ingredients. The continuous functional equation identifies the unique affine candidate (Section 2). Floor discretisation then reduces the identity to a finite counting inequality $N(r)<r\beta+s-1$, where $N(r)$ counts certain fractional parts of multiples of $1/\!\sqrt{r}$ (Proposition 9). A parity argument on $t=r-\lfloor\sqrt{r}\rfloor^{2}$ establishes this inequality for all non-square $r$ (Section 3).
@@ 1.4 The triple-nested family
The identity (3) selects a unique value of $d$ (namely $d=\sqrt{2}/2$), but the weaker identity (5) admits a continuous family.
@@
Let $c=1/\!\sqrt{2}$ and I=[2/2, 2(21)]I=[$\sqrt${2}/2,\;2($\sqrt${2}-1)]. For every $d\in I$, the sequence defined by $a(n)=\lfloor cn+d\rfloor$ satisfies (5) for all $n\geq 1$. The interval $I$ is sharp. For $d\notin I$ the identity fails at infinitely many $n$.
The endpoints of $I$ are the two Beatty parameters associated with the $\sqrt{2}$-Wythoff game (Fraenkel [4]).
@@ 1.5 Outline
Sections 23 prove the existence results. Section 2 establishes the exact continuous solution. Section 3 reduces the discrete identity to a counting inequality via sawtooth analysis, then closes it by a parity argument for all $r$ that are not even perfect squares. Section 4 proves the Beatty interval for the triple-nested equation (Theorem 3). The remaining sections discuss structural phenomena specific to $r=2$: Section 5 discusses the Sturmian structure and bifurcation behaviour in the case $r=2$, and Section 6 studies a conjectural defect morphism at the boundary of the Beatty interval.
@@ 2 The continuous equation
In the discrete equation (1), $f(n)=\sum a(n{-}j)$ denotes the sliding-window sum and the identity reads $a\circ f=\mathrm{id}$. The continuous analogue replaces $a$ by a smooth function $\varphi$:
@@
The affine function
$\varphi(x)=\frac{x}{\sqrt{r}}+\frac{\sqrt{r}-1}{2}$ (7)
satisfies (6), as one checks by substitution.
@@
Affine solutions. If $\varphi(x)=cx+d$ then
@@
Rigidity. Let $\varphi$ be any continuous nondecreasing solution and set $g(x)=\sum_{j=0}^{r-1}\varphi(x{-}j)$. Then $g$ is continuous and nondecreasing, and $\varphi(g(x))=x$ for all $x$, so $\varphi$ is surjective and $g$ is injective. Since $g$ is continuous, nondecreasing, and injective, it is strictly increasing. If $\varphi$ had a flat interval $[u,v]$ with $\varphi\equiv\lambda$ there, then $g(\lambda)\in[u,v]$ but $g(x)<u$ for $x<\lambda$ and $g(x)>v$ for $x>\lambda$, contradicting continuity of $g$ at $\lambda$. So $\varphi$ is strictly increasing, hence a homeomorphism, and $g=\varphi^{-1}$. ∎
Remark 5.
The discrete Beatty solution $a(n)=\lfloor n/\!\sqrt{r}+\sqrt{r}/2\rfloor$ is the floor-discretisation of $\varphi$: the shift $d_{\mathrm{Beatty}}=\sqrt{r}/2$ equals $d_{\mathrm{cont}}+1/2=(\sqrt{r}-1)/2+1/2$, the universal +1/2+1/2 correcting for the average of the fractional part.
@@ 3 The universal identity
3.1 The discretisation residual
By Theorem 4, the continuous identity $\varphi(\sum\varphi(x{-}j))=x$ is exact. The only obstruction to the discrete identity $a(S(n))=n$ is the rounding error introduced by the floor. Define
@@ 3.2 Counting reformulation
The local maximum of the sawtooth just before the jump at $\theta=\{j_{0}c\}$ is
@@
and the local minimum just after the same jump is $M(j_{0})-1$ (using $\{-mc\}=1-\{mc\}$ for $m\geq 1$). Figure 3 shows the sawtooth function over a full period for $r=2,3,5$.
Figure 3: The sawtooth $\Phi_{r}(\theta)=\sum_{j=0}^{r-1}\{\theta-jc\}$ for $r=2,3,5$ with $c=1/\!\sqrt{r}$, over one period $\theta\in[0,1)$.
@@
By Proposition 6 and the density of $\{cn+\alpha/2\}$ in $[0,1)$, the identity holds iff $(r{-}\alpha)/2<\Phi(\theta)\leq(r{+}\alpha)/2$ for all $\theta$. By Lemma 8, $\sup\Phi+\inf\Phi=r$, so the two bounds are equivalent: it suffices to show $\sup\Phi<(r{+}\alpha)/2$ (equality excluded by irrationality of $c$).
@@
Since each step $\Delta_{k}$ lies in $\{\beta{-}1,\,\beta\}$, the partial sum $P(j)=j\beta-C(j)$ where $C(j)=\#\{1\leq k\leq j:\{kc\}<\beta\}$. The sequence $\{kc\}$ is a Sturmian rotation, and the bounded remainder property of irrational rotations gives $|C(j)-j\beta|<1$ for all $j$ (this is the balance property of Sturmian words, see [7, Chapter 2]). Hence $|P(j)|<1$ for every $j$. By Lemma 8, $\max P+\min P=P(r{-}1)$, so $\max P-\min P<2$.
@@ 3.3 The parity argument
Since $\beta=t/(s+\alpha)$ and $s+\alpha\in(2s,2s{+}1)$, we get $t/(2s{+}1)<\beta<t/(2s)$. Multiplying by $s$: $st/(2s{+}1)<s\beta<t/2$. Since st/(2s+1)(t1)/2=(2st+1)/$(2$s+1))>0st/(2s{+}1)-(t{-}1)/2=(2s{-}t{+}1)/$(2$s{+}1))>0 (as $t\leq 2s$), we obtain $(t{-}1)/2<s\beta<t/2$, whence $\lfloor s\beta\rfloor=\lfloor(t{-}1)/2\rfloor$. ∎
@@
We now prove Theorem 2.
Proof of Theorem 2.
By Proposition 9, it suffices to show $N(r)<r\beta+s-1$. By Lemmas 10 and $1$, we split by the parity of $t$.
Case t=2u+1t=2u{+}1 odd. Then $\lfloor s\beta\rfloor=u$, and substituting (14) into Lemma 11 gives $N(r)=2\sum_{m=1}^{u}\lceil m/\beta\rceil$. Since $\beta$ is irrational, each $\lceil m/\beta\rceil<m/\beta+1$, so
@@
Case $t=2u$ even. Then $\lfloor s\beta\rfloor=u{-}1$, and the same substitution gives $N(r)=s{+}1+2\sum_{m=1}^{u-1}\lceil m/\beta\rceil$. Using $\lceil m/\beta\rceil<m/\beta+1$ again:
@@ 3.4 Perfect squares
Case $s$ odd (s=2q+1s=2q{+}1). Then $d=s/2=q+1/2\notin\mathbb{Z}$, and $a(n)=q+\lfloor(n{+}q)/s\rfloor$. For $n\geq s^{2}$ the full window is active, so
@@
As $u$ ranges over $\{0,\ldots,s{-}1\}$, the values n+qun{+}q{-}u hit each residue class modulo $s$ exactly once, so by Hermite’s identity ($\sum_{v=0}^{n-1}\lfloor x+v/n\rfloor=\lfloor nx\rfloor$ for all $x\in\mathbb{R}$, see [6, §3.5])
@@
Case $s$ even ($s=2m$). Then $d=m\in\mathbb{Z}$ and $a(n)=m+\lfloor n/s\rfloor$. Applying Hermite’s identity in the same way gives S(n)=sns(s2)/2S(n)=sn-s(s{-}2)/2. Then $a(S(n))=m+\lfloor n{-}(s{-}2)/2\rfloor=n+1$, since $(s{-}2)/2=m{-}1\in\mathbb{Z}$. ∎
@@ 3.5 Coexistence of solutions
For each $r$ (excluding $r=(2m)^{2}$), the Golomb equation (1) of order $r$ admits at least two different monotone solutions: the greedy [2], which is $r$-regular with oscillating ratio $a(n)/n$, and the Beatty, with $a(n)/n\to 1/\!\sqrt{r}$. Both are fixed points of the implicit equation $a\circ f=\mathrm{id}$, sustained by different self-consistent address structures.
@@ 4 The triple-nested equation
We now compose the Golomb equation (1) of order $r$ with $a$, obtaining the weaker identity $a(a(f(n)))=a(n)$. This reveals a continuous family of Beatty solutions parametrised by a shift $d$ ranging over an explicit interval, with connections to Wythoff games and combinatorics on words.
4.1 The continuous perspective

### 21 Apr 2026 v3

@@ Abstract
The almost Golomb equation of order $r$ is the implicit functional equation
@@
for nondecreasing sequences of positive integers with $a(1)=1$. Its earliest solution, the almost Golomb sequence of order $r$, is $r$-regular in the sense of Allouche and Shallit and has oscillating ratio $a(n)/n$. We prove that for every $r\geq 2$ that is not an even perfect square, the equation admits a second monotone solution given by an inhomogeneous Beatty sequence of slope $1/\!\sqrt{r}$.
Composing the equation with $a$ leads to a triple-nested identity which admits a continuous one-parameter family of inhomogeneous Beatty solutions, parametrised by a shift $d$ ranging over an explicit interval. We determine these intervals sharply for $r=2$ and $r=3$, each proved by a local regime analysis combined with equidistribution of an irrational orbit. The endpoints of these intervals sit naturally inside the Pell–Ostrowski framework of Fokkink, and the defect set at the upper endpoint for $r=2$ is characterised as the return-time set of an irrational rotation to an explicit interval.
@@
Keywords. Beatty sequence, almost Golomb sequence, functional equation, Sturmian word, Pell tower, Ostrowski numeration.
@@ 1.2 The greedy solution
The greedy (earliest) solution of (1) is studied in detail in the companion paper [2]. It is unique, $r$-regular, and has explicit $r$-adic denesting formulas. For $r=2$ (A394217), it begins
@@ 1.3 A second solution
The two solutions agree on $[1,11]$ and diverge at $n=12$, where the greedy repeats the value $8$ (run of length $2$), while the Beatty moves to $9$ (run of length $1$). This is possible because the constraint at $n=12$ applies at position $f(12)=a(12)+a(11)$, which is $16$ for the greedy and $17$ for the Beatty. Both sequences thus satisfy the equation at $n=12$, but with the constraint applied at different positions, allowing divergence of the two solutions from that point on.
@@
The residual $a_{\mathrm{B}}(n)-n/\!\sqrt{2}$ is bounded and equals $\{(n{+}1)/\!\sqrt{2}\}\in(-1,0]$, whereas $a_{\mathrm{G}}(n)-n/\!\sqrt{2}$ grows without bound because $a_{\mathrm{G}}(n)/n$ oscillates in $[2/3,3/4]$ rather than converging to $1/\!\sqrt{2}$.
@@
When $r=s^{2}$ with $s$ even, the left-hand side equals $n+1$ instead of $n$. When $s$ is odd the identity holds as in the non-square case (Theorem 13).
The proof combines three ingredients. The continuous functional equation identifies the unique affine candidate (Section 2). Floor discretisation then reduces the identity to a finite counting inequality on fractional parts of multiples of $1/\!\sqrt{r}$ (Proposition 9). A parity argument on $t=r-\lfloor\sqrt{r}\rfloor^{2}$ establishes this inequality for all non-square $r$ (Section 3).
@@ 1.4 The triple-nested family
Write the inhomogeneous Beatty sequence of slope $c=1/\!\sqrt{2}$ and shift $d$ as
$a(n)=\lfloor cn+d\rfloor.$
The canonical solution (2) corresponds to $d=\sqrt{2}/2$, which is the unique shift for which the strong identity (3) holds. The weaker triple-nested identity (5), by contrast, admits a continuous family of shifts. The paper determines this interval sharply for $r=2$, and the analogous interval for $r=3$.
@@
Let $c=1/\!\sqrt{2}$ and I=[2/2, 2(21)]I=[$\sqrt${2}/2,\;2($\sqrt${2}-1)]. The sequence defined by $a(n)=\lfloor cn+d\rfloor$ satisfies (5) for all $n\geq 2$ if and only if $d\in I$.
The analogous theorem for $r=3$, stated and proved in Section 4.6, gives
$a(a(a(n)+a(n{-}1)+a(n{-}2)))=a(n)\quad\text{for all }n\geq 3\iff d\in I_{3}=\Bigl[\frac{\sqrt{3}+3}{6},\;\frac{5\sqrt{3}-3}{6}\Bigr].$
Both intervals are proved sharp by the same method, namely a local analysis of the four combinatorial regimes induced by the Sturmian rotation combined with equidistribution of the irrational orbit. Convergents of $c$ play no role in the proof.
The right endpoint $d=2(\sqrt{2}{-}1)$ for $r=2$ admits a simple algebraic form. At this boundary the strong identity fails on a sparse defect set. We prove (Theorem 24) that this defect set is the return-time set of the Sturmian rotation $\theta\mapsto\{\theta+c\}$ to the interval [1c,1/2][1{-}c,1/2], which explains every failure by absorption into a repeat of $a_{\mathrm{B}}$. The three-distance theorem then yields three gap values $\{3,4,7\}$. Numerical evidence suggests that the gap sequence is generated by an explicit primitive substitution with Perron eigenvalue $1+\sqrt{2}$, formulated as Conjecture 26. A full proof via Rauzy induction on three letters is left for future work.
The equation, the interval $I$, and the defect structure all sit naturally inside the Pell–Ostrowski framework of Fokkink [4], which generalises the Wythoff array of Conway and Ryba to the recurrence $X_{n+1}=2X_{n}+X_{n-1}$. We make this connection explicit in Section 4.5, and note that the family $r=s^{2}+1$ for $s\geq 1$ corresponds to Fokkink’s Ostrowski towers of parameter $2s$.
@@ 1.5 Outline
Sections 23 prove the existence results. Section 2 establishes the exact continuous solution. Section 3 reduces the discrete identity to a counting inequality via sawtooth analysis, then closes it by a parity argument for all $r$ that are not even perfect squares. Section 4 is the rigorous core of the paper. It proves the Beatty intervals for the triple-nested equation at $r=2$ (Theorem 3) and at $r=3$ (Theorem 21), and describes their Pell–Ostrowski interpretation. The remaining sections discuss structural phenomena specific to $r=2$. Section 5 discusses the Sturmian structure and bifurcation behaviour. Section 6 proves that the defect set at the upper endpoint of the Beatty interval is the return-time set of a Sturmian rotation to an explicit sub-interval (Theorem 24), computes its density and the three gap values, and formulates as a conjecture the primitive substitution on $\{3,4,7\}$ whose Perron eigenvalue is the silver ratio $1+\sqrt{2}$.
@@ 2 The continuous equation
In the discrete equation (1), $f(n)=\sum a(n{-}j)$ denotes the sliding-window sum and the identity reads $a\circ f=\mathrm{id}$. The continuous analogue replaces $a$ by a continuous nondecreasing function $\varphi$:
@@
A natural candidate is the affine function
$\varphi(x)=\frac{x}{\sqrt{r}}+\frac{\sqrt{r}-1}{2}.$ (7)
@@
Suppose first that $\varphi(x)=cx+d$ is affine. Then
@@
We now show that every continuous nondecreasing solution is a homeomorphism satisfying the inverse equation. Let $\varphi$ be any such solution and set $g(x)=\sum_{j=0}^{r-1}\varphi(x{-}j)$. Then $g$ is continuous and nondecreasing, and $\varphi(g(x))=x$ for all $x$, so $\varphi$ is surjective and $g$ is injective. Since $g$ is continuous, nondecreasing, and injective, it is strictly increasing. If $\varphi$ had a flat interval $[u,v]$ with $\varphi\equiv\lambda$ there, then $g(\lambda)\in[u,v]$ but $g(x)<u$ for $x<\lambda$ and $g(x)>v$ for $x>\lambda$, contradicting continuity of $g$ at $\lambda$. So $\varphi$ is strictly increasing, hence a homeomorphism, and $g=\varphi^{-1}$. ∎
@@ 3 The general identity
The goal of the section is the following, announced as Theorem 2 in the introduction.
Theorem 5.
For every $r\geq 2$ that is not an even perfect square, $a(S(n))=n$ for all $n\geq r$. When $r=s^{2}$ with $s$ even, $a(S(n))=n+1$ for all $n\geq r$. When $r=s^{2}$ with $s$ odd, the identity $a(S(n))=n$ holds as in the non-square case.
The proof occupies Sections 3.1 through 3.4. It proceeds in three steps. Section 3.1 reduces the identity $a(S(n))=n$ to a uniform bound on the sawtooth function $\Phi$ (Proposition 6). Sections 3.23.3 reduce that bound to a counting inequality on fractional parts of multiples of $c$, and close the inequality by a parity argument on $t=r-\lfloor\sqrt{r}\rfloor^{2}$. Section 3.4 treats the perfect-square cases separately via Hermite’s identity.
3.1 Reduction to a bound on $\Phi$
By Theorem 4, the continuous identity $\varphi(\sum\varphi(x{-}j))=x$ is exact. The only obstruction to the discrete identity $a(S(n))=n$ is the rounding error introduced by the floor. Define the sawtooth function
@@ 3.2 Counting reformulation
The local maximum of the sawtooth just before the jump at $\theta=\{j_{0}c\}$, where $j_{0}\in\{1,\ldots,r{-}1\}$ indexes the jump points, is
@@
and the local minimum just after the same jump is $M(j_{0})-1$ (using $\{-mc\}=1-\{mc\}$ for $m\geq 1$).
@@
By Proposition 6, the identity $a(S(n))=n$ for all $n\geq r$ is equivalent to $(r{-}\alpha)/2<\Phi(n)\leq(r{+}\alpha)/2$ for all such $n$. Since $\theta=\{n/\alpha+\alpha/2\}$ is equidistributed in $[0,1)$ by Weyl’s theorem (using that $c=1/\alpha$ is irrational), and $\Phi(n)$ depends on $n$ only through $\theta$, the bound extends to the sawtooth $\Phi\colon[0,1)\to\mathbb{R}$ viewed as a function of the orbit parameter $\theta$. By Lemma 8, $\sup\Phi+\inf\Phi=r$, so the two bounds are equivalent: it suffices to show $\sup\Phi<(r{+}\alpha)/2$ (equality excluded by irrationality of $c$).
@@
Since each step $\Delta_{k}$ lies in $\{\beta{-}1,\,\beta\}$, the partial sum $P(j)=j\beta-C(j)$ where $C(j)=\#\{1\leq k\leq j:\{kc\}<\beta\}$. The orbit $(\{kc\})_{k\geq 1}$ is a Sturmian rotation, and the bounded remainder property of irrational rotations gives $|C(j)-j\beta|<1$ for all $j$ (this is the balance property of Sturmian words, see [7, Chapter 2]). Hence $|P(j)|<1$ for every $j$. By Lemma 8, $\max P+\min P=P(r{-}1)$, so $\max P-\min P<2$.
@@
Remark 10.
The value $N(r)=0$ observed at $r=2$ and $r=5$ in Table 1 is not an anomaly. One has $N(r)=0$ precisely when $r=s^{2}+1$ for some $s\geq 1$ (sequence A002522), the values for which $\sqrt{r}$ has continued-fraction expansion $s;\overline{2s}]$ of minimal period one. In that case $t=1$ and the counting inequality $N(r)<r\beta+s-1$ reduces to $r\beta+s-1>0$, obvious for $s\geq 1$. The next such values are $r=10,\,17,\,26,\,37,\ldots$
@@ 3.3 The parity argument
Since $\beta=t/(s+\alpha)$ and $s+\alpha\in(2s,2s{+}1)$, we get
$\frac{t}{2s{+}1}<\beta<\frac{t}{2s}.$
Multiplying by $s$:
$\frac{st}{2s{+}1}<s\beta<\frac{t}{2}.$
Since
$\frac{st}{2s{+}1}-\frac{t{-}1}{2}=\frac{2s{-}t{+}1}{2(2s{+}1)}>0$
(as $t\leq 2s$), we obtain $(t{-}1)/2<s\beta<t/2$, whence $\lfloor s\beta\rfloor=\lfloor(t{-}1)/2\rfloor$. ∎
@@
We now prove the non-square part of Theorem 5.
Proof of Theorem 5, non-square case.
By Proposition 9, it suffices to show $N(r)<r\beta+s-1$. By Lemmas 11 and 12, we split by the parity of $t$.
Suppose first that t=2u+1t=2u{+}1 is odd. Then $\lfloor s\beta\rfloor=u$, and substituting (14) into Lemma 12 gives $N(r)=2\sum_{m=1}^{u}\lceil m/\beta\rceil$. Since $\beta$ is irrational, each $\lceil m/\beta\rceil<m/\beta+1$, so
@@
Suppose now that $t=2u$ is even. Then $\lfloor s\beta\rfloor=u{-}1$, and the same substitution gives $N(r)=s{+}1+2\sum_{m=1}^{u-1}\lceil m/\beta\rceil$. Using $\lceil m/\beta\rceil<m/\beta+1$ again:
@@ 3.4 Perfect squares
Suppose first that s=2q+1s=2q{+}1 is odd. Then $d=s/2=q+1/2\notin\mathbb{Z}$, and $a(n)=q+\lfloor(n{+}q)/s\rfloor$. For $n\geq s^{2}$ the full window is active, so
@@
As $u$ ranges over $\{0,\ldots,s{-}1\}$, the values n+qun{+}q{-}u hit each residue class modulo $s$ exactly once, so by Hermite’s identity ($\sum_{v=0}^{n-1}\lfloor x+v/n\rfloor=\lfloor nx\rfloor$ for all $x\in\mathbb{R}$)
@@
Suppose now that $s=2m$ is even. Then $d=m\in\mathbb{Z}$ and $a(n)=m+\lfloor n/s\rfloor$. Applying Hermite’s identity in the same way gives S(n)=sns(s2)/2S(n)=sn-s(s{-}2)/2. Then $a(S(n))=m+\lfloor n{-}(s{-}2)/2\rfloor=n+1$, since $(s{-}2)/2=m{-}1\in\mathbb{Z}$. ∎
@@ 3.5 Coexistence of solutions
For each $r$ (excluding $r=(2m)^{2}$), the Golomb equation (1) of order $r$ admits at least two different monotone solutions. The greedy [2] is $r$-regular with oscillating ratio $a(n)/n$. The Beatty has $a(n)/n\to 1/\!\sqrt{r}$. Both are fixed points of the implicit equation $a\circ f=\mathrm{id}$, sustained by different self-consistent address structures.
@@ 4 The triple-nested equation
We now compose the Golomb equation (1) for $r=2$ with $a$, obtaining the weaker identity $a(a(f(n)))=a(n)$. This reveals a continuous family of Beatty solutions parametrised by a shift $d$ ranging over an explicit interval. The goal of this section is the following, announced as Theorem 3 in the introduction.
Theorem 16.
Let $c=1/\!\sqrt{2}$, f(n)=a(n)+a(n1)f(n)=a(n)+a(n{-}1), and $a(n)=\lfloor cn+d\rfloor$. The triple-nested identity $a(a(f(n)))=a(n)$ holds for all $n\geq 2$ if and only if
$d\in I=\bigl[\sqrt{2}/2,\;2(\sqrt{2}-1)\bigr].$
The proof occupies Sections 4.14.4. Necessity is handled by two failure lemmas that locate, in each of the three residue classes of $\theta_{n}=\{cn+d\}$, a sub-interval where the identity (5) fails (Section 4.3). Orbit density, via Weyl equidistribution, then produces infinitely many failures. Sufficiency is a case analysis of the same residue classes (Section 4.4). Section 4.5 then places the interval II and its endpoints inside Fokkink’s Pell–Ostrowski framework. Section 4.6 proves the analogous result for $r=3$ by the same method.
4.1 The continuous equation